OpenStudy (anonymous):

Verify the identity using co-function identities for sine and cosine and basic identites cot(pi/2 -x)=tanx

OpenStudy (anonymous):

haha and what are you supposed to do with this? :>

OpenStudy (anonymous):

I forgot to put what I was supposed to do, I just fixed it.

OpenStudy (anonymous):

doesn't seem so bad, once you realise what cot really is... ^^

OpenStudy (anonymous):

well, i got to (cos(pi/2-x))/sin(pi/2-x)=sinx/cosx but i dont know what to do next

OpenStudy (anonymous):

sin/cos is tan, plain and simple. basically, you've already done it :D

OpenStudy (anonymous):

then I cosx/-sinx=sinx/cosx

OpenStudy (anonymous):

wait what? lol sinx/cosx is already equal to tanx, it's a basic identity ^^

OpenStudy (anonymous):

i know, but, cos(pi/2-x)/(sin pi/2 -x) = cospi/2 + cos(-x) / sin pi/2 sin(-x)

OpenStudy (anonymous):

seems like you're complicating things... cos(pi/2 - x) = sin(x) and sin(pi/2 - x) = cos(x) these should suffice lawl

OpenStudy (anonymous):

then cos pi/2= 0 cos -x= cosx sin pi/2=1 sin(-x)= -sinx

OpenStudy (anonymous):

why do you get sin(-x) ?

OpenStudy (anonymous):

That makes more sense, thank you!

OpenStudy (anonymous):

oh ok... :>

OpenStudy (anonymous):

I know what i did wrong, cosx=sin(pi/2 -x) sinx=cos(pi/2 -x)

OpenStudy (anonymous):

like I said, lol. it's ok, as long as you saw your mistakes.. that's how we learn ^.^